thee more than 4 times a number equal to 9 less than twice the number

Found 2 solutions by stanbon, rfer:
Answer by stanbon(75887)

three more than 4 times a number is equal to 9 less than twice the number
----------------
Let the number be "x":
Equation:
4x+3 = 2x-9
2x = -12
x = -6
==================
Cheers,
Stan H.

Answer by rfer(16322)

4x+3=2x-9
2x=-12
x=-6
-----------------
4*-6+3=2*-6-9
-24+3=-12-9
-21=-21